# Why Vision Language Models Struggle with Visual Arithmetic? Towards Enhanced Chart and Geometry Understanding

### v2

@@ Abstract
Kung-Hsiang Huang1    Can Qin1    Haoyi Qiu2 Philippe Laban1    Shafiq Joty1    Caiming Xiong1    Chien-Sheng Wu1 1Salesforce AI Research    2UCLA 1{kh.huang, cqin, sjoty, wu.jason}@salesforce.com
@@ 1 Introduction
Despite these advancements, current VLMs exhibit noticeable deficiencies in performing fundamental visual arithmetic: these models struggle with seemingly simple tasks like accurately counting objects, comparing lengths, assessing angles, and evaluating relative sizes or areas Rahmanzadehgervi_2024_blind; wang2024vdlm; huang2024frompixels; ullman2024illusion; wei2024slow. These shortcomings are particularly evident in complex tasks such as chart understanding huang-etal-2024-lvlms and geometric problem-solving gao2023gllava.
@@ 2.2 Probing Analysis
Table 2: Accuracy (%) of fine-tuned LLaVA-v1.5-7B with different queries on the test set of each probing task. We conduct experiments by freezing the vision encoder and only fine-tune the LLM decoder for binary classification.
@@
Based on the above findings and analyses, we conclude that while visual representations from VLMs do encompass sufficient information for visual arithmetic, it cannot be effectively decoded without further fine-tuning, which may in turn affect their zero-shot performance on downstream tasks like chart understanding.

### v3

@@ Abstract
Kung-Hsiang Huang1    Can Qin1    Haoyi Qiu2 Philippe Laban1    Shafiq Joty1    Caiming Xiong1    Chien-Sheng Wu1 1Salesforce AI Research    2UCLA 1{kh.huang, cqin, sjoty, cxiong, wu.jason}@salesforce.com
@@ 1 Introduction
Despite these advancements, current VLMs exhibit noticeable deficiencies in performing fundamental visual arithmetic: these models struggle with seemingly simple tasks like accurately counting objects, comparing lengths, assessing angles, and evaluating relative sizes or areas Rahmanzadehgervi_2024_blind; wang2024vdlm; huang2024frompixels; ullman2024illusion; wei2024slow; kamoi2024visonlyqa. These shortcomings are particularly evident in complex tasks such as chart understanding huang-etal-2024-lvlms and geometric problem-solving gao2023gllava.
@@ 2.2 Probing Analysis
Table 2: Accuracy (%) of fine-tuned LLaVA-v1.5-7B with different queries on the test set of each probing task. We conduct experiments by freezing the vision encoder and only fine-tune the LLM decoder for binary classification. The top row displays the zero-shot performance.
@@
2.3 Visual Representation Inspection
While the analysis conducted in Section 2.2 reveals important findings, one may argue that the fine-tuning success might be caused by the fine-tuned model exploiting patterns in binary tasks. Motivated by the visual representation analysis of rudman2025forgotten, we conduct additional experiments focusing on visual representations’ ability to differentiate isolated geometric properties. We generated datasets of images, each containing a single angle (sampled in 10-degree increments from 10° to 100°) or a single line (sampled in 0.1 length increments from 0.1 to 1.0). We created 100 images per class, resulting in 1,000 images each for the angle and line datasets. We then extracted visual embeddings for these images using the pre-trained CLIP and SigLIP vision encoders. To assess whether embeddings for different angles or lengths occupy distinct regions in the feature space (a prerequisite for successful differentiation or classification), we performed t-SNE visualization vandermaaten08tsne to understand cluster separation. The results are shown in Figure 2.
We have the following observations. First, there are some clear clusters for both CLIP and SigLIP embeddings, particularly for smaller angles and shorter lines. However, the separation is less distinct for longer lines and larger angles, especially with CLIP embeddings. Second, the separation for angles is more obvious than that for lines, aligning with our findings in Table 3 that Length Comparison is more challenging than Angle Comparison. Third, SigLIP embeddings demonstrate superior clustering for both angles and lines. This is also consistent with our Table 3 results where SigLIP outperforms CLIP across the proposed probing tasks.
Based on these observations and our findings in Section 2.2, we conclude that while visual representations from VLMs do encode information for visual arithmetic, it cannot be effectively decoded without further fine-tuning, which may in turn affect their zero-shot performance on downstream tasks like chart understanding and that fine-tuning success is not merely exploiting patterns in a binary task.
(a) Angle clustering by CLIP
(b) Line clustering by CLIP
(c) Angle clustering by SigLIP
(d) Line clustering by SigLIP
Figure 2: t-SNE visualizations of clustering by CLIP and SigLIP for angles and lines.